OpenStudy (anonymous):

can so one help me to change: z1=4i; z2=-6+6i to polar form???? can so one help me to change: z1=4i; z2=-6+6i to polar form???? @Mathematics

hero (hero):

Ugh...Hate these

OpenStudy (anonymous):

Yes: \[z_{1} = 4i = 4 (\cos(\frac{\pi}{2}) + i\sin(\frac{\pi}{2}))\] \[z_{2} = -6+6i = 6\sqrt{2} (\cos(\frac{3\pi}{2}) + i\sin(\frac{3\pi}{2}))\]

OpenStudy (anonymous):

hihihi i tried to make it already but it wasn't come out right!!:(

OpenStudy (anonymous):

how about if i need to find the rectangular form????

hero (hero):

I need to review my trig

OpenStudy (anonymous):

ok.... can u help me???? because i tried already n it came out not right!!:(..

hero (hero):

I could help you, but not today. I would really need to review fully.

OpenStudy (anonymous):

ok thanks:)....

hero (hero):

Do you have skype?

OpenStudy (anonymous):

yes i do.... but i cant log in my skype account now...

OpenStudy (phi):